# An NV- center in magnesium oxide as a spin qubit for hybrid quantum technologies

### v1

@@ II.4 Vibronic analysis
According to the character table of the $C_{4v}$ point group (see SI Table III), the $E$ electronic states can couple linearly to the $a_{1}$, $a_{2}$, $b_{1}$, $b_{2}$ phonon modes, and to $e$ phonon modes via higher (even) order coupling. However, only the $b_{1}$, $b_{2}$, and $e$ phonon modes can reduce the symmetry of the defect. Given that the $C_{4v}$ symmetry is reduced to $C_{1}$ in the optimized geometry of the first excited ${}^{3}E$ state, $e$-phonon modes are expected to be predominantly coupled to the electronic states ($b_{1}$ and $b_{2}$ modes would have reduced the symmetry to $C_{2}$ and $C_{2v}$, respectively). To verify that indeed the $e$ phonon modes are responsible for the observed symmetry-breaking in the first excited ${}^{3}E$ state, we calculated the Mg, O, and N displacements between the average structure of the four equivalent minima (shown in Fig. 4a) and the optimized geometry of the excited state (shown in Fig. 3a). The average structure is symmetrical and the displacement between the average geometry and that of the relaxed excited state geometry may be used to understand the predominant symmetry-breaking mode. We found that the symmetry-breaking displacements are predominantly due to the $e$-phonon mode (see SI Section III and insets of Fig.4b).
@@
To quantify the coupling between the various states, we fit an adiabatic potential energy surface (APES) to the first excited state CCD, which yields an effective phonon frequency of the $e$-phonon mode and the vibronic coupling constants. Specifically, we solve a $(E+E)\otimes e$ Hamiltonian, which accounts for Jahn-Teller (JT) coupling within each ${}^{3}E$ state and pseudo-Jahn Teller (pJT) coupling between two excited ${}^{3}E$ states. We used an adiabatic-to-diabatic transformation of the potential energy matrix and built an effective Hamiltonian that includes coupling constants for JT and pJT effects [58, 59]. We then fit the eigenvalues of the Hamiltonian to the TDDFT-DDH energy values along the CCD of the first excited state to obtain the parameters of the energy matrix. The details of the derivation are given in the SI Section IV. The APES of the first excited ${}^{3}E$ state along the $e$-phonon mode is shown in Fig.4b, from which we obtained the effective phonon frequency of the $e$-phonon mode to be 25 meV, and the vibronic coupling parameters to be 20-500 meV. We also found that the pJT effect is more pronounced than the JT effect: including the JT coupling parameters did not reduce the fitting error of the APES (see SI Section IV for more details). The pJT stabilization energy is large ($\sim$5 eV). These vibronic coupling parameters are far larger than those of the NV- in diamond (in which vibronic coupling is around 120 meV [60]), and the effective phonon frequency is far lower than that of the NV- in diamond (effective phonon frequency of the latter is around 60 meV [60]). The large pJT stabilization energy and low effective phonon frequencies satisfy the condition for strong vibronic coupling, given as [55]
@@ II.5 Engineering strategies to reduce vibronic coupling and increase DWF
To check if this reduction leads to a reduced HRF and increased DWF, we followed the same strategy as described in Section C. We constructed a one-dimensional CCD along the direction connecting the optimized geometries in the ground and excited state using TDDFT-DDH and fit a quadratic function around the points in proximity of the respective minima. In this way, we obtained effective phonon frequencies of the ground and excited states under the different strain conditions, which we used to calculate the HRF. We find that strain decreases the HRF from 44.49 (unstrained case) to 41.36 (1% strained case) and 36.23 (4% strained case). This results in an increase in the DWF by 4 and 7 orders of magnitude respectively; however, the ZPL would still be undetectable, as the DWF of the unstrained system is extremely low ($\approx$ 10-20).
@@ III Discussion
In this work, we used the ADAQ high-throughput screening framework, along with hybrid DFT and TDDFT methods and QDET, to discover and characterize a promising NV--like spin defect in MgO. The oxide host satisfies all the desirable criteria listed in the seminal paper by Weber et al. [65]: it has a wide band gap, small spin-orbit coupling, is available as high-quality thin films and bulk crystals, and both Mg and O have naturally occurring isotopes of zero nuclear spin and thus can be isotopically purified. Additionally, the recent study by Kanai et al. [5] predicted a single spin in MgO to have a coherence time ($T_{2}$) of 0.60 ms. This is slightly lower than the $T_{2}$ in prototypical hosts like diamond (0.89 ms) or SiC (1.1 ms) (primarily due to the larger concentration of nuclear spin isotopes of Mg). However, MgO could still be a promising host due to other advantages, particularly the ability to be grown as a high quality thin film via multiple routes [17, 18], and its use in a wide range of applications. These include MgO tunnel barriers in spintronics devices [66], MgO buffer layers in superconducting [20] and ferroelectric [19] devices, and MgO heterogeneous catalyst support materials [67].
@@
In addition to the favorable properties of the host, many properties of the NV- defect in MgO are promising as well. In particular, its spin triplet ground state provides a two-level system ($m_{s}$ = 0 and $m_{s}$ = $\pm$1), and the large ZFS can enable isolation of the two sublevels. The large ZFS and generally low hyperfine interaction parameters could be beneficial to increase the coherence time of the defect compared to the estimate obtained by considering a single electron spin in MgO. The singlet states between the ${}^{3}A_{2}$ ground and first ${}^{3}E$ excited state, as determined using TDDFT-DDH and QDET (see Fig. 3b), can act as shelving states, thereby enabling optical initialization and spin sublevel-dependent photo-luminescence. The efficiency of the initialization will depend on the radiative and non-radiative transition rates, which will be the subject of a future study. The presence of singlet shelving states can also enable additional readout modes, for example spin-to-photocurrent (or spin-to-charge) readout. Another favorable feature of the NV- defect in MgO is its emission in the visible red-infrared regime, making the experimental detection of this defect relatively straightforward. Moreover, it is unlikely that the emitted light will be reabsorbed, as one may expect oxygen vacancies to be the additional most likely defect to be present in the oxide, and its photoexcitation and photoionization energies are around 5 eV [68]. The broad emission line-width of the NV- center may allow filtering of any emitted light from underlying substrates, while still retaining a substantial portion of emitted light from the defect which can be detected. Importantly, MgO can be doped with nitrogen, as demonstrated in previous experimental studies [28, 29, 30, 31, 32, 33]. The hyperfine interaction parameters of the NV- center are different from those of other nitrogen-related defects, e.g. the nitrogen interstitial and nitrogen substitutional on the oxygen site [38], which would make the NV- center experimentally distinguishable.
@@
The properties of the NV- spin defect computed here and the wide technological applicability of MgO thin films point at applications that could integrate ‘classical’ spintronics and ferroelectrics with quantum information science. For example, classical information, e.g. magnetization direction in a magnetic tunnel junction (MTJ), with MgO as the tunnel barrier, can be mapped onto the spin defect, as it has been recently proposed by Jansen and Yuasa [69] in the case of a quantum dot. The NV- defect could also be used in conjunction with MgO MTJ-based nanomagnets to realize single qubit gates for quantum computing [70]. Similarly, it may be interesting to explore mapping of the polarization state of a ferroelectric thin film to control qubit interactions in MgO buffer layers [71]. The properties of the NV- in MgO could also make it a viable quantum sensor. In fact, flexible and wearable spintronic devices based on MgO-barrier MTJs have been recently developed [72]: combining the NV- spin defect with this device could enable wearable quantum sensing.
@@
A key challenge with the promising defect identified in our study is the strong vibronic coupling, which leads to a large $\Delta Q$ and consequently large HRF and low DWF, leading to an undetectable ZPL. This likely makes the NV- center in MgO unsuitable for quantum communication; however, quantum sensing and computing could be viable applications. Our study revealed that strain could help increase the DWF, however, one would need to go to extremely large strains to achieve a detectable ZPL. Note that the DWF can be increased via the formation of point-planar defect complexes, like in SiC nanowires, where complexes formed between point defects and stacking faults were found to have high brightness and DWF $\approx$ 50%. [57]. It is important to note that the large $\Delta Q$ and resulting low DWF found in our study at zero temperature represents the worst-case scenario, since all our analysis was done starting from a highly symmetric ground state structure. However, the NV- defect in MgO has multiple nearly-degenerate local minima with lower symmetry. Given that zero-point vibrational energy in MgO is non-negligible (it leads to lattice constant expansion by 0.5 % and bandgap renormalization by 6 % [73, 74]), it is likely that the zero-point vibrational energy will lead to symmetry breaking, which will reduce $\Delta Q$ and increase the DWF. Indeed, if we consider structures with broken symmetry in the ground state (as reported in our ADAQ search, see SI Table VIII), we find a decrease of the $\Delta Q$.
@@
Importantly, our study was able to shed light onto the reasons behind the strong vibronic coupling found for the NV- in MgO. Our APES analysis showed that the strong vibronic coupling is due to the large pJT stabilization energy in combination with the low effective phonon frequency along the symmetry-breaking $e$-phonon modes. This knowledge gives insight into design principles that could be considered while exploring or screening new spin defects, especially for enhancing optical addressability and/or enabling quantum communication applications. In particular, the strength of the vibronic coupling due to the pJT effect increases when the energy separation between excited states decreases[55]. This behavior is often correlated with the energy separation between the single-particle Kohn-Sham defect levels. Thus, to lower the vibronic coupling (therefore lowering $\Delta Q$ and HRF), the localized mid-gap states (both occupied and unoccupied) must be well-separated from each other, which is what we observe in substitutional complexes such as the $\mathrm{Al_{Mg}-Al_{O}}$ defect in MgO (compare Fig. 5b with Fig. 2b for the NV-). Moreover, the HRF increases when the charge distribution in the ground and excited state differ substantially in localization[54], which is the case for the NV- defect in MgO (where the excitation is from the O $2p$ defect states around the magnesium vacancy to the $\pi$-orbital around the N-O bond). Instead, in VV0 in SiC, NV- in diamond, and even the $\mathrm{Al_{Mg}-Al_{O}}$ defect in MgO, the ground- and excited-state orbitals are localized in the same region of space. Thus, we expect spin defects in oxides comprising substitutional defects (including $d$- and $f$-dopants) to have lower vibronic coupling, leading to lower $\Delta Q$ and increased DWF. Moreover, our study shows that $\Delta Q$, the separation between Kohn-Sham levels, and the localization of the electronic states involved in the excitation are good descriptors for screening or designing spin defects.
@@
In sum, the computational protocol utilized in our study, combining high-throughput techniques and accurate electronic structure methods, helped identify and characterize a promising NV- defect in MgO. This new spin defect in a technologically mature material may be integrated into spintronic/ferroelectric applications to realize multifunctional electronics, including spintronic quantum logic gates and sensors. Overall, our results also contribute to the understanding of oxides as hosts for spin defects.
@@ IV.1 ADAQ
We used the ADAQ [21, 22] software package and the high-throughput toolkit [23] to create and screen defects in MgO, and the automatic screening workflow is presented in Ref. 21.
The MgO lattice parameter was optimized with the Perdew, Burke, and Erzenerhof (PBE) functional  [75], and we obtained 4.25 Å. We used supercells with 512 atoms for all the defect calculations, and the Lany-Zunger charge correction [76] with a dielectric constant $\epsilon_{0}$ = 9.65. In ADAQ, the calculations were performed with the Vienna Ab initio Simulation Package (VASP) [77, 78], gamma compiled version 5.4.4, which uses the projector augmented wave (PAW) [79, 80] pseudopotential (folder dated 2015-09-21). We used a plane-wave energy cutoff of 600 eV and kinetic energy cutoff of 900 eV for the augmentation charges, a total energy criterion of $10^{-4}$ eV, and a structural minimization criterion of $5\times 10^{-3}$ eV.
@@ IV.2 Electronic structure calculations
Once the NV-- center was identified through the ADAQ procedure, we computed its ground state properties at the DFT level using Quantum ESPRESSO [81, 82, 83]. We used scalar-relativistic SG15 ONCV pseudopotentials (v. 1.2) [84, 85] and the dielectric-dependent hybrid (DDH) [41] exchange-correlation functional, with a plane-wave cutoff of 80 Ry. The inverse of the experimental macroscopic dielectric constant $\epsilon_{\infty}$ was used to determine the fraction of exact exchange $\alpha$ ($\epsilon_{\infty}$ = 2.96 [42, 41], yielding $\alpha$ = $\epsilon_{\infty}^{-1}$ = 0.34) . Given the computational cost of hybrid functionals, we used supercells with 216 atoms and we fixed the lattice constant to the experimental value extrapolated to 0 K, including the correction for zero-point anharmonic expansion effects [41] ($a_{0}$ = 4.19 Å). The supercell Brillouin zone was sampled with the $\Gamma$ point. Forces were converged to 0.01 eV/Å, and all calculations were spin-polarized.
@@
where $E^{f}[D^{q}]$ is the total energy of the supercell containing the defect $D$ in charge state $q$, $E_{tot}[bulk]$ is the total energy of the pristine supercell of the same size, $n_{i}$ is the number of atoms of species $i$ added to ($n_{i}>0$) or removed from ($n_{i}<0$) the supercell to create the defect, $\mu_{i}$ is the chemical potential of species $i$, $\varepsilon_{f}$ is the Fermi energy relative to the valence band maximum $E_{VBM}$, and $E_{corr}$ is a charge-dependent term that corrects for artificial electrostatic interactions arising due to the finite size of the supercell and was calculated using the Freysoldt-Neugebauer-Van de Walle correction scheme [87] with the static dielectric constant for MgO $\epsilon_{0}=9.8$ taken from experiment [88]. Note that a comparison between the Lany-Zunger and Freysoldt-Neugebauer-Van de Walle correction schemes in the case of CaO [6] yielded very similar and fully consistent results, and we expect the same for MgO.

### v2

@@ II.4 Vibronic analysis
According to the character table of the $C_{4v}$ point group (see Supplementary Table 5), the $E$ electronic states can couple linearly to the $a_{1}$, $a_{2}$, $b_{1}$, $b_{2}$ phonon modes, and to $e$ phonon modes via higher (even) order coupling. However, only the $b_{1}$, $b_{2}$, and $e$ phonon modes can reduce the symmetry of the defect. Given that the $C_{4v}$ symmetry is reduced to $C_{1}$ in the optimized geometry of the first excited ${}^{3}E$ state, $e$-phonon modes are expected to be predominantly coupled to the electronic states ($b_{1}$ and $b_{2}$ modes would have reduced the symmetry to $C_{2}$ and $C_{2v}$, respectively). To verify that indeed the $e$ phonon modes are responsible for the observed symmetry-breaking in the first excited ${}^{3}E$ state, we calculated the Mg, O, and N displacements between the average structure of the four equivalent minima (shown in Fig. 4a) and the optimized geometry of the excited state (shown in Fig. 3a). The average structure is symmetrical and the displacement between the average geometry and that of the relaxed excited state geometry may be used to understand the predominant symmetry-breaking mode. We found that the symmetry-breaking displacements are predominantly due to the $e$-phonon mode (see Supplementary Note 6 and insets of Fig.4b).
@@
To quantify the coupling between the various states, we fit an adiabatic potential energy surface (APES) to the first excited state CCD, which yields an effective phonon frequency of the $e$-phonon mode and the vibronic coupling constants. Specifically, we solve a $(E+E)\otimes e$ Hamiltonian, which accounts for Jahn-Teller (JT) coupling within each ${}^{3}E$ state and pseudo-Jahn Teller (pJT) coupling between two excited ${}^{3}E$ states. We used an adiabatic-to-diabatic transformation of the potential energy matrix and built an effective Hamiltonian that includes coupling constants for JT and pJT effects [67, 68]. We then fit the eigenvalues of the Hamiltonian to the TDDFT-DDH energy values along the CCD of the first excited state to obtain the parameters of the energy matrix. The details of the derivation are given in Supplementary Note 7. The APES of the first excited ${}^{3}E$ state along the $e$-phonon mode is shown in Fig.4b, from which we obtained the effective phonon frequency of the $e$-phonon mode to be 25 meV, and the vibronic coupling parameters to be 20-500 meV. We also found that the pJT effect is more pronounced than the JT effect: including the JT coupling parameters did not reduce the fitting error of the APES (see Supplementary Note 7 for more details). The pJT stabilization energy is large ($\sim$5 eV). These vibronic coupling parameters are far larger than those of the NV- in diamond (in which vibronic coupling is around 120 meV [69]), and the effective phonon frequency is far lower than that of the NV- in diamond (effective phonon frequency of the latter is around 60 meV [69]). The large pJT stabilization energy and low effective phonon frequencies satisfy the condition for strong vibronic coupling, given as [59]
@@
We end this section by briefly mentioning that the strong vibronic coupling could indicate a strong spin-phonon coupling, which might lead to a reduction in the spin relaxation time $T_{1}$. While a spin-phonon coupling analysis is beyond the scope of this study, we did calculate the phonon modes for the NV- defect in the ground state, as discussed in Supplementary Note 8. We find that the modes are largely localized near the NV- defect and have energies that generally exceed its ZFS. Thus, the spin relaxation time $T_{1}$ might be limited by slower two-phonon processes as opposed to fast one-phonon direct and Orbach relaxation mechanisms. Having predominantly two-phonon processes would be desirable for an optimal value of the defect’s $T_{1}$.
@@ II.5 Engineering strategies to reduce vibronic coupling and increase DWF
To check if this reduction in $\Delta Q$ leads to a reduced HRF and increased DWF, we followed the same strategy as described in Section C. We constructed a one-dimensional CCD along the direction connecting the optimized geometries in the ground and excited state using TDDFT-DDH and fit a quadratic function around the points in proximity of the respective minima. In this way, we obtained effective phonon frequencies of the ground and excited states under the different strain conditions, which we used to calculate the HRF. We find that strain decreases the HRF from 44.49 (unstrained case) to 41.36 (1% strained case) and 36.23 (4% strained case). This results in an increase in the DWF by 4 and 7 orders of magnitude respectively; however, unfortunately the ZPL would still be undetectable, as the DWF of the unstrained system is extremely low ($\approx$ 10-20).
@@ III Discussion
In this work, we used the ADAQ high-throughput screening framework, along with hybrid DFT and TDDFT methods and QDET, to discover and characterize a promising NV- spin defect in MgO. The oxide host satisfies all the desirable criteria listed in the seminal paper by Weber et al. [74]: it has a wide band gap, small spin-orbit coupling, it is available as high-quality thin films and bulk crystals, and both Mg and O have naturally occurring isotopes of zero nuclear spin and thus can be isotopically purified. Additionally, the recent study by Kanai et al. [5] predicted a single spin in MgO to have a coherence time ($T_{2}$) of 0.60 ms. This is slightly lower than the $T_{2}$ in prototypical hosts like diamond (0.89 ms) or SiC (1.1 ms) (primarily due to the larger concentration of nuclear spin isotopes of Mg). However, MgO could still be a promising host due to other advantages, particularly the ability to be grown as a high quality thin film via multiple routes [17, 18], and its use in a wide range of applications. These include MgO tunnel barriers in spintronics devices [75], MgO buffer layers in superconducting [21] and ferroelectric [19, 20] devices, and MgO heterogeneous catalyst support materials [76].
@@
In addition to the favorable properties of the host, many properties of the NV- defect in MgO are promising as well. In particular, its spin triplet ground state provides a two-level system ($m_{s}$ = 0 and $m_{s}$ = $\pm$1), and the large ZFS can enable isolation of the two sublevels. The large ZFS and generally low hyperfine interaction parameters could be beneficial to increase the coherence time of the defect compared to the estimate obtained by considering a single electron spin in MgO. The singlet states between the ${}^{3}A_{2}$ ground and first ${}^{3}E$ excited state, as determined using TDDFT-DDH and QDET (see Fig. 3b), can act as shelving states, thereby enabling optical initialization and spin sublevel-dependent photo-luminescence. The efficiency of the initialization will depend on the radiative and non-radiative transition rates, which will be the subject of a future study. However, we note that the absence of crossing (or near-degeneracy) between the potential energy surfaces of the singlet and triplet states is an important indicator of the possibility of optical initialization and readout. The presence of singlet shelving states can also enable additional readout modes, for example spin-to-photocurrent (or spin-to-charge) readout. Another favorable feature of the NV- defect in MgO is its emission in the visible red-infrared regime, making the experimental detection of this defect relatively straightforward. It is unlikely that the emitted light will be reabsorbed, as one may expect oxygen vacancies to be the additional most likely defect to be present in the oxide, and its photoexcitation and photoionization energies are around 5 eV [77]. The broad emission line-width of the NV- center may allow filtering of any emitted light from underlying substrates, while still retaining a substantial portion of emitted light from the defect which can be detected. Moreover, recent work on Er3+-doped MgO [66] has demonstrated the integration of MgO with silicon nanophotonic cavities to achieve a cavity enhancement of over three orders of magnitude. This points to the promising possibility of using cavity enhancement of the optical transitions in the NV- center as well. Importantly, MgO can be doped with nitrogen, as demonstrated in previous experimental studies [29, 30, 31, 32, 33, 34]. The hyperfine interaction parameters of the NV- center are different from those of other nitrogen-related defects, e.g. the nitrogen interstitial and nitrogen substitutional on the oxygen site [39], indicating that the NV- center should be experimentally distinguishable.
@@
The properties of the NV- spin defect computed here and the wide technological applicability of MgO thin films point at applications that could integrate ‘classical’ spintronics and ferroelectrics with quantum information science. For example, classical information, e.g. magnetization direction in a magnetic tunnel junction (MTJ), with MgO as the tunnel barrier, can be mapped onto the spin defect, as it has been recently proposed by Jansen and Yuasa [78] in the case of a quantum dot. The NV- defect could also be used in conjunction with MgO MTJ-based nanomagnets to realize single qubit gates for quantum computing [79]. Similarly, it may be interesting to explore mapping of the polarization state of a ferroelectric thin film to control qubit interactions in MgO buffer layers [80, 81, 82]. Further, the properties of the NV- in MgO could make it a viable quantum sensor. In fact, flexible and wearable spintronic devices based on MgO-barrier MTJs have been recently developed [83]: combining the NV- spin defect with these devices could enable wearable quantum sensing.
@@
A key challenge with the promising defect identified in our study is the strong vibronic coupling, which leads to a large $\Delta Q$ and consequently large HRF and low DWF, leading to an undetectable ZPL. This likely makes the NV- center in MgO unsuitable for quantum communication; however, quantum sensing and computing could be viable applications. Our study revealed that strain could help increase the DWF, however, one would need to go to extremely large strains to achieve a detectable ZPL. Note that the DWF can be increased via the formation of point-planar defect complexes, like in SiC nanowires, where complexes formed between point defects and stacking faults were found to have high brightness and DWF $\approx$ 50%. [61]. It is important to note that the large $\Delta Q$ and resulting low DWF found in our study at zero temperature represents the worst-case scenario, since all our analysis was done starting from a highly symmetric ground state structure. However, the NV- defect in MgO has multiple nearly-degenerate local minima with lower symmetry. Given that zero-point vibrational energy in MgO is non-negligible (it leads to a lattice constant expansion by 0.5 % and bandgap renormalization by 6 % [84, 85]), it is likely that the zero-point vibrational energy will lead to symmetry breaking, which will reduce $\Delta Q$ and increase the DWF. Indeed, if we consider structures with broken symmetry in the ground state (as reported in our ADAQ search, see Supplementary Table 1), we find a decrease of the $\Delta Q$.
@@
Importantly, our study was able to shed light onto the reasons behind the strong vibronic coupling found for the NV- in MgO. Our APES analysis showed that the strong vibronic coupling is due to the large pJT stabilization energy in combination with the low effective phonon frequency along the symmetry-breaking $e$-phonon modes. This finding gives insight into design principles that could be considered while exploring or screening new spin defects, especially for enhancing optical addressability and/or enabling quantum communication applications. In particular, the strength of the vibronic coupling due to the pJT effect increases when the energy separation between excited states decreases[59]. This behavior is often correlated with the energy separation between the single-particle Kohn-Sham defect levels. Thus, to lower the vibronic coupling (therefore lowering $\Delta Q$ and HRF), the localized mid-gap states (both occupied and unoccupied) must be well-separated from each other, which is what we observe in substitutional complexes such as the $\mathrm{Al_{Mg}-Al_{O}}$ defect in MgO (compare Fig. 5b with Fig. 2b for the NV-). Moreover, the HRF increases when the charge distribution in the ground and excited state differ substantially in localization[58], which is the case for the NV- defect in MgO (where the excitation is from the O $2p$ defect states around the magnesium vacancy to the $\pi$-orbital around the N-O bond). Instead, in VV0 in SiC, NV- in diamond, and even in the $\mathrm{Al_{Mg}-Al_{O}}$ defect in MgO, the ground- and excited-state orbitals are localized in the same region of space. Thus, we expect spin defects in oxides comprising substitutional defects (including $d$- and $f$-dopants) to have lower vibronic coupling, leading to lower $\Delta Q$ and increased DWF. Moreover, our study shows that $\Delta Q$, the separation between Kohn-Sham levels, and the localization of the electronic states involved in the excitation are good descriptors for screening or designing spin defects.
@@
In sum, the computational protocol utilized in our study, combining high-throughput techniques and accurate electronic structure methods, helped identify and characterize a promising NV- defect in MgO. This new spin defect in a technologically mature material may be integrated into spintronic/ferroelectric applications to realize multifunctional electronics, including spintronic quantum logic gates and sensors. Overall, our results on the vibronic coupling of the NV- defect in MgO also contribute to the understanding of oxides as hosts for spin defects.
@@ IV.1 ADAQ
We used the ADAQ [22, 23] software package and the high-throughput toolkit [24] to create and screen defects in MgO, and the automatic screening workflow is presented in Ref. 22.
The MgO lattice parameter was optimized with the Perdew, Burke, and Erzenerhof (PBE) functional  [86], and we obtained 4.25 Å. We used supercells with 512 atoms for all the defect calculations, and the Lany-Zunger charge correction [87] with a dielectric constant $\epsilon_{0}$ = 9.65. In ADAQ, the calculations were performed with the Vienna Ab initio Simulation Package (VASP) [88, 89], gamma compiled version 5.4.4, which uses the projector augmented wave (PAW) [90, 91] pseudopotential (folder dated 2015-09-21). We used a plane-wave energy cutoff of 600 eV and kinetic energy cutoff of 900 eV for the augmentation charges, a total energy criterion of $10^{-4}$ eV, and a structural minimization criterion of $5\times 10^{-3}$ eV. Calculations were done at the $\Gamma$ point. All the 2917 defects screened can be found in the ADAQ database online.
@@ IV.2 Electronic structure calculations
Once the NV-- center was identified through the ADAQ procedure, we computed its ground state properties at the DFT level using Quantum ESPRESSO [92, 93, 94]. We used scalar-relativistic SG15 ONCV pseudopotentials (v. 1.2) [95, 96] and the dielectric-dependent hybrid (DDH) [42] exchange-correlation functional, with a plane-wave cutoff of 80 Ry. Convergence tests for the plane-wave cutoff energy are provided in Supplementary Note 10. The inverse of the experimental macroscopic dielectric constant $\epsilon_{\infty}$ was used to determine the fraction of exact exchange $\alpha$ ($\epsilon_{\infty}$ = 2.96 [43, 42], yielding $\alpha$ = $\epsilon_{\infty}^{-1}$ = 0.34) . Given the computational cost of hybrid functionals, we used supercells with 216 atoms and we fixed the lattice constant to the experimental value extrapolated to 0 K, including the correction for zero-point anharmonic expansion effects [42] ($a_{0}$ = 4.19 Å). Note that this closely matches the DDH optimized lattice parameter (4.195 Å) as obtained from an equation-of-state fit (see Supplementary Note 11). The supercell Brillouin zone was sampled with the $\Gamma$ point. Forces were converged to 0.01 eV/Å, and all calculations were spin-polarized.
@@
where $E^{f}[D^{q}]$ is the total energy of the supercell containing the defect $D$ in charge state $q$, $E_{tot}[bulk]$ is the total energy of the pristine supercell of the same size, $n_{i}$ is the number of atoms of species $i$ added to ($n_{i}>0$) or removed from ($n_{i}<0$) the supercell to create the defect, $\mu_{i}$ is the chemical potential of species $i$, $\varepsilon_{f}$ is the Fermi energy relative to the valence band maximum $E_{VBM}$, and $E_{corr}$ is a charge-dependent term that corrects for artificial electrostatic interactions arising due to the finite size of the supercell and was calculated using the Freysoldt-Neugebauer-Van de Walle correction scheme [98] with the static dielectric constant for MgO $\epsilon_{0}=9.8$ taken from experiment [99]. Note that a comparison between the Lany-Zunger and Freysoldt-Neugebauer-Van de Walle correction schemes in the case of CaO [6] yielded very similar and fully consistent results, and we expect the same for MgO. The chemical potentials of oxygen, nitrogen, and magnesium were calculated as $\mu_{O}=\frac{1}{2}E_{DFT,O_{2}}$, $\mu_{N}=\frac{1}{2}E_{DFT,N_{2}}$, $\mu_{Mg}=E_{DFT,MgO}-\mu_{O}$. We used the HSE functional [100] to obtain $E_{DFT,O_{2}}$ and $E_{DFT,N_{2}}$, as the binding energy and the bond length of the respective molecules calculated using the HSE functional matched more closely with experiments. Additional details can be found in Supplementary Note 12. A comparison of the defect formation energies computed at the PBE (the functional used in the ADAQ screening procedure) and DDH levels of theory is given in the Supplementary Information as well (Supplementary Note 13).